Grade 12Physical Chemistry

A sample of air is saturated with benzene (vapour pressure =100mm hg at298k)at 298k 750 mm hg pressure . If it is isothermally compressed to one third of its initial volume ,the final pressure of the system is

Komal kumari
8 Years agoGrade 12

1 Answer

Abhiraj Pathak
ApprovedApproved Tutor Answer8 Years ago
Pressure of air = 750-100 = 650mmHgApplying Boyles law,Pnew = 650*3 = 1950mmHg1950 mmHg is the pressure of air only after compression.Since vapor pressure of benzene will remain same at isothermal condition,Psystem = 1950+100 =2050mmHgKindly confirm and approve